OpenStudy (anonymous):

Please help me use Factor Theorem to determine if the given binomial is a factor of the given polynomial function: f(x) = 4x^4 - 20x^3 + 33x^2 - 22x + 5; x - 5/2

OpenStudy (anonymous):

(A) f(-5/2) = 0; yes, the binomial is a factor of the polynomial (B) f(5/2)=0 ; yes, the binomial is a factor of the polynomial (C) f(-5/2) =/= 0; the binomial is not a factor of the polynomial. (D) f(5/2) =/= 0; the binomial is not a factor of the polynomial

OpenStudy (anonymous):

only B and D make sense, check those

OpenStudy (anonymous):

when i say "check" i mean "cheat" you are going to have to use a calculator anyways it turns out \(\frac{5}{2}\) is a zero http://www.wolframalpha.com/input/?i=+4x^4+-+20x^3+%2B+33x^2+-+22x+%2B+5

OpenStudy (anonymous):

So my answer is B? @satellite73